OpenStudy (anonymous):

Find the equation of the tangent line to the graph of the function f(x)=(3x-2)/x+3 at the point (2, 0.8)

OpenStudy (anonymous):

f'(x)= 5/(x+3)^2 ===> f'(2)=1/5 ===> y=(1/5)x+0.4

OpenStudy (anonymous):

where did you get your 5

OpenStudy (anonymous):

i got 11

OpenStudy (ybarrap):

f(x)= [(3x-2)/x] + 3 f'(x) = [x(3) - (3x - 2)(1)] / x^2 , derivative using quotient rule f'(2) = (6 - 6 + 2)/4, slope of f at x = 2 = (2)/4 = 1/2 equation of tangent line: ytan = f'(2)x + b ytan(2) = .8 = 1(2)/2 + b b = .8 - 1 = 8/10 - 1 = (8 - 10)/10 = -1/5 Ans. ytan = x/2 - 1/5